Scores on a Statistics Final Exam are normally distributed. Using a​ 95% confidence​ level, find the sample size necessary to predict the population mean within 3 points. Assume the population standard deviation is 8 points.
Answer #1

Solution :

Given that,

standard deviation =s =  \sigma =8

Margin of error = E = 3

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

sample size = n = [Z\alpha/2* \sigma / E] 2

n = ( 1.96* 8 /3 )2

n =27

Sample size = n =27